# Johannes Matzat        Aiko Schmeißer

### v3

@@ Abstract
Labor unions’ greatest potential for political influence likely arises from their direct connection to millions of individuals at the workplace. There, they may mobilize and change the ideological positions of both unionizing workers and their non-unionizing management. In this paper, we analyze the workplace-level impact of unionization on workers’ and managers’ political campaign contributions in the United States from 1980 to 2016. To do so, we link establishment-level union election data with transaction-level campaign contributions to federal and local candidates. Using a difference-in-differences design, validated through regression discontinuity tests and a novel instrumental variable approach, we find that unionization leads to a leftward shift of campaign contributions. Unionization increases the support for Democrats relative to Republicans not only among workers but also among managers, suggesting that it does not exacerbate political cleavages between the two groups. Our findings are not driven by compositional changes of the workforce and are stronger in establishments where, after union certification, the employer and the union agree on a collective bargaining agreement.
@@ 1 Introduction
Labor unions have long been credited with shaping key welfare and labor market policies, such as the 8-hour workday, minimum wages, safety standards, and sickness and family leave (e.g., Biden, 2021; King, 1965; Obama, 2010). While economists typically attribute unions’ influence to their impact on work contracts via collective bargaining and study their effects on wages and other benefits (Card, 1996; DiNardo and Lee, 2004; Farber et al., 2021; Frandsen, 2021; Knepper, 2020), other scholars have seen unions as essential drivers of social and political change (Baumgartner and Leech, 1998; Dahl, 2005; Lipset, 1960). Through supporting Democratic candidates financially and lobbying for labor-friendly policies, unions are often viewed as one of the few vehicles that gives political voice to workers and enhances their representation in U.S. politics (Burns et al., 2000; Dark, 2001; Rosenfeld, 2014; Schlozman, 2015). Yet, whether unions can bring about lasting change in welfare states and policies depends on their ability to mobilize and change the political preferences of workers and the broader public.
@@
Unions’ greatest political leverage likely arises from their connection to more than 14 million union members and their colleagues at unionized workplaces. [^1] Unions draw on significant resources to finance their political outreach. In 2010, labor unions in the United States employed over 3,000 full-time political workers and spent 700 million USD on political activities, a figure that rose to 1.8 billion USD in 2020 (WSJ, 2012; NILRR, 2021). After family and friends, the workplace is one of the most important arenas for social interaction and political discussion (Hertel-Fernandez, 2020), making it a particularly influential space for unions. By providing political information and training as well as facilitating communication networks between members, unions can mobilize workers and affect their ideological positions. Still, the aggregate political influence of unions at the workplace is far from clear. Even if unions successfully rally unionized workers around their political positions, it is unclear whether they can also persuade the firm’s management. Heightened tension between workers and managers, who represent the owners’ interests, could lead to adverse responses to labor issues. Any backlash in the political behavior of this powerful out-group may prevent unions from achieving their political agenda.
@@
In this paper, we examine the influence of labor unions on the political participation and political ideologies of employees in the United States. We combine establishment-level data on 6,063 union elections with transaction-level data on 357,436 campaign contributions made to federal and local candidates from 1980 to 2016. In the campaign contribution data, we observe the employer, occupation, and address of individual donors, which allows us to match donors from different occupations with the union election results of their employing establishments. To estimate the causal effects of unionization, we employ a difference-in-differences (DiD) framework, comparing campaign contributions of employees in establishments that voted for unionization with those that voted against it (tests of the underlying parallel trends assumption and alternative sources of exogenous variation are described below). We assess the political effects of unions by examining two key outcomes: political mobilization, measured by changes in employees’ total contribution amounts, and ideological shifts, captured by changes in the party composition of candidates they donate to. Linking these outcomes to union elections at the establishment level offers various new opportunities for studying the political influence of unions.
@@
To start with, it enables us to analyze the political effects of unions at the workplace, where unions directly engage with employees and where not only members but also non-members can be affected by unionization. Exploiting the occupational information in the campaign contribution data, we can differentiate the political responses of workers and managers, allowing us to explore within-firm dynamics that have not been studied before. We first ask how workplace unionization alters the political behavior of workers. As Kerrissey and Schofer (2013) argue, unions provide their members with political capital – they inform, engage, and mobilize members. Unions spend substantial resources on outreach and political education to inform members about issues relevant for their working conditions (Ahlquist and Levi, 2013; Iversen and Soskice, 2015; Macdonald, 2021). Moreover, employee gatherings, voting for union officers, participation in hiring halls, and joint strike activities can improve communication networks between workers and create social experiences that turn them into more engaged citizens (Lindvall, 2013; McAdam et al., 2001; Terriquez, 2011).
@@
Through these mechanisms, workplace unionization can increase union members’ participation and support for Democratic candidates, which is consistent with our results. Our DiD estimates show that total campaign contributions from workers rise by 11% in response to unionization. This effect only shows up in the cycle of the union election, suggesting a short-term political mobilization of workers in response to a successful union campaign. Most importantly, when we examine the party composition of contributions, we find that unionization increases the percentage difference in donations from workers to Democrats versus Republicans by 12 percentage points in the six years following a union election. This result indicates a lasting shift in workers’ ideological positions towards the political left.
@@
Focusing solely on union members would overlook a critical out-group: the firm’s management, whose responses can either amplify or counteract the political influence of unions. On one hand, unionization may foster greater understanding of worker issues. Unions give workers a collective voice, presenting their preferences on a more equal footing with management (Freeman and Medoff, 1979, 1984). The establishment of clear bargaining rules may improve both the quantity and quality of communication between managers and workers (Verma, 2005). According to contact theory, this increase in cooperative interactions can enhance perspective-taking and, ultimately, lead to an alignment of ideological positions (e.g., Allport, 1954). [^2] Ash et al. (2019) find that granting workers greater authority through entitlements in collective bargaining agreements reduces labor conflicts, as reflected by a decrease in the frequency and intensity of strikes following negative wage shocks. On the other hand, labor unions may provoke a backlash from management. Representing the interests of firm owners, managers are typically profoundly hostile to unionization. [^3] In the run-up to union elections, employers frequently hire anti-union law firms and consultants, try to delay the election process, hold meetings in which employees are obligated to listen to the anti-unionization arguments, and – although legally restricted – threaten employees with dismissals and establishment closures (Flanagan, 2007; Freeman and Kleiner, 1990a; Kleiner, 2001; Logan, 2002; Schmitt and Zipperer, 2009). The group threat hypothesis posits that the increased bargaining power of workers may be perceived by managers as undermining their own status, power, and economic interests (e.g., Sherif et al., 1961). When unionization heightens the salience of distributional conflicts and reinforces worker-manager identities, it may increase political polarization, as individuals increasingly adopt the stereotypes associated with these identities (Bonomi et al., 2021).
@@
Thus, it is unclear ex ante whether labor unions can persuade managers to align with workers’ political positions or whether they reinforce the management’s opposition. By exploiting occupational information in the campaign contribution data, we can directly estimate the response of managers. Our results show a leftward shift in campaign contributions not only for workers but also for managers: unionization increases the relative difference in managers’ donations to Democratic versus Republican candidates by 20 percentage points, without affecting their total spending. These results suggest that, instead of increasing tensions, unionization fosters a convergence in ideological positions between workers and their management.
@@
Combining an establishment-level political outcome with variation in unionization after union elections also provides us with plausible identification strategies to identify the causal impact of unionization on employees’ political behavior. Given that we only consider establishments with union elections, i.e., where workers have shown an interest in unionization, our sample can be expected to be more similar than a random sample of establishments. Within this sample, we compare campaign contributions between establishments where workers voted for and against unionization by estimating a stacked DiD model. The stacked DiD approach addresses weighting issues that arise under staggered treatment timing and heterogeneous treatment effects (Goodman-Bacon, 2021). [^4] We also check the robustness of our results to employing alternative DiD estimators for settings with staggered treatment timing (Borusyak et al., forthcoming; Callaway and Sant’Anna, 2021), which replicate our main results from the stacked DiD model.
@@
The DiD design relies on the assumption that, absent unionization, campaign contributions in losing establishments would have followed the same trend as in winning establishments. The plausibility of that assumption is validated by complementing the DiD framework with tests originating from a regression discontinuity design (RDD) and with a novel instrumental variables (IV) approach. First, we test whether changes in outcomes are correlated with the pro-union vote share among establishments that lost the union election. Since the treatment status changes discontinuously at the 50% threshold, there should be no differential trends across establishments with different vote shares below 50%. Indeed, we do not find any evidence for differential changes across different vote shares, which helps us to rule out the possibility that any sizeable confounding factors correlated with the pro-union vote share and the timing of the election drive the results. Second, we restrict the sample to establishments with increasingly close elections, which are more likely to follow similar trends in contribution patterns. Our results are robust to a wide range of vote share bandwidths around the 50% cutoff, even when focusing on elections decided by only a 5-10% margin. Finally, we complement the DiD with arguably exogenous variation in union support, exploiting random shocks to the salience of workplace safety that are triggered by unexpected spikes in sector-level fatal work accidents shortly before the union election. The DiD-IV results confirm our main findings.
@@
The establishment-level effects of unionization on workers’ and managers’ campaign contributions could be explained by changes in the composition of the workforce and management. To differentiate between compositional and individual-level effects, we use donor identifiers and track each donor’s campaign contributions over time. We develop two specifications. First, we examine compositional changes by estimating the effect on pre-election contribution patterns of donors who donated and were matched to an establishment after the election. We find no sizeable effect, indicating that the pre-existing political preferences of separated and newly hired employees do not differ across unionizing and non-unionizing establishments. Second, we study individual-level effects by focusing on incumbent employees who were employed and donated at the establishment before the union election. We find a significant leftward shift in donations from individual workers and managers. In sum, these results suggest that labor unions successfully persuade both members and their management to support Democratic candidates.
@@
Our data also allows us to move beyond party preferences by considering candidates’ ideological positions and endorsement by union organizations, as well as employees’ support of interest groups. We document considerable within-party variation in the effects on contributions to different candidates. Liberal candidates gain and conservative candidates lose, while moderate candidates are not significantly impacted on average. This suggests that our findings are not only driven by an increased signal of Democratic over Republican party affiliation but reflect shifts between candidates with clearly distinguishable ideological positions. Moreover, we differentiate candidates by whether or not they are supported by unions’ Political Action Committees (PACs). The results indicate that the partisan shift goes beyond union-endorsed candidates, reflecting a more general alignment with Democratic candidates. In addition, we show that our results extend to contributions from employees to PACs: unions mobilize workers to increase donations to union and membership PACs, while managers reduce contributions to corporate PACs. The increased support for labor and civil society interest groups from workers and the reduced support for business interest groups from managers match with the observed ideological realignment in their contributions to candidates. At the same time, we find no significant increase in managers’ donations to union PACs, challenging an alternative strategic explanation under which managers merely signal goodwill toward workers and the union.
@@
Third, we shed new light on the spread of political preferences at work by combining establishment-level union election data with an individual-level political outcome. Existing research on the political impact of unions has focused either on individual union members and their households (e.g., Freeman, 2003) or on aggregate outcomes comprising the whole county or state population (e.g., Feigenbaum et al., 2018). By focusing on the unionizing workplace, we are the first to consider within-firm dynamics, particularly the reaction of management –- an out-group that is likely indirectly affected by unionization and a key actor in political influence. Thereby, we expand on studies of politics at the workplace (Colonnelli et al., 2022; Fos et al., 2023), specifically those examining spillovers in political donations between managers and workers (Babenko et al., 2020; Stuckatz, 2022) and the political effects of intergroup contact at work (Andersson and Dehdari, 2021).
@@ 2.1 Unionizing through NLRB Elections
Since 1935, the National Labor Relations Act (NLRA) gives most private-sector workers in the U.S. the right to organize in unions and take collective action, such as bargaining and strikes. Collective bargaining between unions and employers takes place at the establishment level. Traditionally, workers unionize through a secret ballot election at their establishment that is administered by the National Labor Relations Board (NLRB). [^7] While union elections are the primary means by which private-sector workers gain union representation, there are alternative procedures for unionization. First, employers may voluntarily recognize unions without an election through neutrality agreements and “card checks”. These cases are less common, however, since employers generally oppose union organization (Schmitt and Zipperer, 2009). Second, some workers’ bargaining rights are not regulated by the NLRA. For example, the Railway Labor Act determines bargaining rights of airline and railroad workers and several federal, state, and local laws regulate the organization of public-sector employees. The unionization procedure involves three main steps: a petition drive, an election, and certification. [^8] The description of the unionization process follows Frandsen (2021) and Wang and Young (2023).
The organizing drive can be initiated either by the workers at an establishment or by a union organization. The initiator first needs to gather the signatures of at least 30% of workers in the proposed bargaining unit who thereby express a desire for unionization. With these signatures, an election petition is filed to the NLRB. The NLRB decides whether to accept the petition by ascertaining whether workers in the proposed bargaining unit share common interests that can be adequately represented by the union. If the petition is accepted, the NLRB schedules a secret ballot election, which usually takes place at the workplace. The union wins the election if it obtains a strict majority of the votes cast. In case of union victory, the NLRB certifies the union as the sole authorized representative of employees in the bargaining unit.

### v4

@@ Abstract
Labor unions influence economic outcomes not only through bargaining with employers over work contracts but also via political activities that can profoundly shape political systems. In unionized workplaces, they may mobilize and change the ideological positions of both unionizing workers and their non-unionizing management. In this paper, we analyze the workplace-level impact of unionization on workers’ and managers’ political campaign contributions in the United States from 1980 to 2016. To do so, we link establishment-level union election data with transaction-level campaign contributions to federal and local candidates. Using a difference-in-differences design, validated through regression discontinuity tests and a novel instrumental variable approach, we find that unionization leads to a leftward shift of campaign contributions. Unionization increases the support for Democrats relative to Republicans not only among workers but also among managers, suggesting that it fosters political alignment between the two groups. Our findings are not driven by compositional changes of the workforce and are stronger in establishments where, after union certification, the employer and union agree on a collective bargaining agreement.
@@ 1 Introduction
Labor market institutions are typically evaluated by economists based on their ability to solve market failures and improve economic efficiency. Yet, the social welfare implications of labor market institutions also hinge critically on their potential ramifications for the political system (Acemoglu and Robinson, 2013). Labor unions have historically functioned as institutions dedicated to redistributing both economic and political power within society. While a long tradition of research has focused on how unions affect work contracts of their members via collective bargaining (Card, 1996; DiNardo and Lee, 2004; Farber et al., 2021; Frandsen, 2021; Knepper, 2020), unions seek to shape economic outcomes more broadly by engaging in various political activities. They are often viewed as one of the few vehicles that give political voice to workers, counterbalancing the dominant influence of business elites in politics (Acemoglu and Johnson, 2023; Rosenfeld, 2014; Schlozman, 2015). [^1] Unions have been credited with shaping key welfare and labor market policies, such as the 8-hour workday, minimum wages, safety standards, and sickness and family leave (e.g., Biden, 2021; King, 1965; Obama, 2010). They invest significant resources into political activities. In 2010, labor unions in the United States employed over 3,000 full-time political workers and spent 700 million USD on political outreach, a figure that rose to 1.8 billion USD in 2020 (NILRR, 2021; WSJ, 2012).
@@
The greatest political leverage of U.S. labor unions likely stems from their connection with more than 14 million union members and their colleagues in unionized workplaces. After family and friends, the workplace is one of the most important arenas for social interaction and political discussion (Hertel-Fernandez, 2020), making it a particularly influential space for unions. By providing political information and training as well as facilitating communication networks between members, unions can mobilize workers and affect their ideological positions. Nonetheless, the overall political influence of unions in the workplace is far from clear. Even if unions successfully rally unionized workers around their political positions, it is unclear whether they can also persuade the firm’s management. Heightened tensions between workers and managers, who act on behalf of ownership interests, may provoke adverse reactions to labor demands. Any backlash in the political behavior of this powerful out-group may prevent unions from shifting political power toward worker interests.
@@
In this paper, we examine the effect of labor unions on the political participation and political ideologies of workers and managers in the United States. We combine establishment-level data on 6,063 union elections with transaction-level data on 357,436 campaign contributions made to federal and local candidates from 1980 to 2016. In the campaign contribution data, we observe the employer, occupation, and address of individual donors, which allows us to match donors from different occupations with the union election results of their employing establishments. To estimate the causal effects of unionization, we employ a difference-in-differences (DiD) framework, comparing campaign contributions of employees in establishments that voted for unionization with those that voted against it (tests of the underlying parallel trends assumption and alternative sources of exogenous variation are described below). We assess the political effects of unions by examining two key outcomes: political mobilization, measured by changes in employees’ total contribution amounts, and ideological shifts, captured by changes in the party composition of candidates they donate to. Linking these outcomes to union elections at the establishment level offers various new opportunities for studying the political influence of unions.
@@
To start with, our approach enables us to analyze the political effects of unions in the workplace, where both unionizing workers and their non-unionizing management can be affected. Exploiting the occupational information in the campaign contribution data, we can differentiate the political responses of workers and managers, allowing us to explore within-firm dynamics that have not been studied before. We first ask how workplace unionization alters the political behavior of workers. As Kerrissey and Schofer (2013) argue, unions provide their members with political capital – they inform, engage, and mobilize members. Unions dedicate significant resources to political outreach and education, informing members on issues related to their working conditions and typically advocating Democratic Party positions (Ahlquist and Levi, 2013; Iversen and Soskice, 2015; Macdonald, 2021). Moreover, employee gatherings, voting for union officers, participation in hiring halls, and joint strike activities can improve communication networks between workers and create social experiences that turn them into more engaged citizens (Lindvall, 2013; McAdam et al., 2001; Terriquez, 2011). Through these mechanisms, unionization can change the political participation and ideological positions of their members.
@@
Our main results document that successful workplace unionization leads to significant increases in workers’ political donations, benefiting Democratic candidates. We find an increase in total campaign contributions from workers by 11% in response to unionization. This effect only shows up in the cycle of the union election, suggesting a short-term political mobilization of workers in response to a successful union campaign. Most importantly, when we examine the party composition of contributions, we find that unionization increases the percentage difference in donations from workers to Democrats versus Republicans by 12 percentage points in the six years following a union election. This result indicates a lasting shift in workers’ ideological positions towards the political left.
@@
Focusing solely on union members would overlook a critical out-group: the firm’s management, whose responses can either amplify or counteract the political influence of unions. On one hand, unionization may foster greater understanding of worker issues. Unions give workers a collective voice, presenting their preferences on a more equal footing with management (Freeman and Medoff, 1979, 1984). The establishment of clear bargaining rules may improve both the quantity and quality of communication between managers and workers (Verma, 2005). According to contact theory, this increase in cooperative interactions can enhance perspective-taking and, ultimately, lead to an alignment of ideological positions (e.g., Allport, 1954). [^2] Ash et al. (2019) find that granting workers greater authority through entitlements in collective bargaining agreements reduces labor conflicts, as seen in fewer and less intense strikes after negative wage shocks. On the other hand, labor unions may provoke a backlash from management. Representing the interests of firm owners, managers are typically profoundly hostile to unionization. [^3] In the run-up to union elections, employers frequently hire anti-union law firms and consultants, try to delay the election process, hold meetings in which employees are obligated to listen to the anti-unionization arguments, and – although legally restricted – threaten employees with dismissals and establishment closures (Flanagan, 2007; Freeman and Kleiner, 1990a; Kleiner, 2001; Logan, 2002; Schmitt and Zipperer, 2009). The group threat hypothesis posits that the increased bargaining power of workers may be perceived by managers as undermining their own status, power, and economic interests (e.g., Sherif et al., 1961). When unionization heightens the salience of distributional conflicts and reinforces worker-manager identities, it may increase political polarization, as individuals increasingly adopt the stereotypes associated with these identities (Bonomi et al., 2021). Thus, it is unclear ex ante whether labor unions can persuade managers to align with workers’ political positions or whether they reinforce the management’s opposition.
@@
Our results notably reveal a leftward shift in campaign contributions not only for workers but also for managers. By exploiting occupational information in the campaign contribution data, we can directly estimate the response of managers. We find that unionization increases the relative difference in managers’ donations to Democratic versus Republican candidates by 20 percentage points, without affecting their total spending. These results suggest that, instead of increasing tensions, unionization fosters a convergence in ideological positions between workers and their management.
@@
Combining establishment-level political outcomes with variation in unionization after union elections also provides plausible identification strategies to identify the causal impact of unionization on employees’ political behavior. Since we only consider establishments with union elections, i.e., where workers have shown an interest in unionization, our sample can be expected to be more similar than a random sample of establishments. Within this sample, we compare campaign contributions between establishments where workers voted for and against unionization by estimating a stacked DiD model. The stacking approach addresses weighting issues that arise under staggered treatment timing and heterogeneous treatment effects (Goodman-Bacon, 2021). [^4] Our results are robust to employing alternative DiD estimators for settings with staggered treatment timing (Borusyak et al., forthcoming; Callaway and Sant’Anna, 2021).
@@
We validate the underlying parallel trends assumption by complementing the DiD framework with tests originating from a regression discontinuity design (RDD) and with a novel instrumental variables (IV) approach. First, we test whether changes in outcomes are correlated with the pro-union vote share among establishments that lost the union election. Since the treatment status changes discontinuously at the 50% threshold, there should be no differential trends across establishments with different vote shares below 50%. Indeed, we do not find any evidence for differential changes across different vote shares, which helps us to rule out the possibility that any sizeable confounding factors correlated with the pro-union vote share and the timing of the election drive the results. Second, we restrict the sample to establishments with increasingly close elections, which are more likely to follow similar trends in contribution patterns. Our results are robust to a wide range of vote share bandwidths around the 50% cutoff, even when focusing on elections decided by only a 5-10% margin. Finally, we complement the DiD with arguably exogenous variation in union support, exploiting random shocks to the salience of workplace safety that are triggered by unexpected spikes in sector-level fatal work accidents shortly before the union election. The DiD-IV results confirm our main findings.
@@
The establishment-level effects of unionization on employees’ campaign contributions could be explained by changes in the contributions of individual employees or by changes in the composition of the employed workforce. To differentiate between individual-level and compositional effects, we use donor identifiers and track each donor’s campaign donations over time. We develop two specifications. First, we study individual-level effects by focusing on incumbent employees who donated and were employed at the establishment before the union election. We find a significant leftward shift in donations from individual workers and managers. Second, we examine compositional changes by estimating the effect on pre-election donation patterns of donors who were matched to an establishment after the election. We find no sizeable effect, indicating that the pre-existing political preferences of separated and newly hired employees do not differ across unionizing and non-unionizing establishments. In sum, these results suggest that labor unions successfully persuade both workers and their management to support Democratic candidates.
@@
Our data also allows us to move beyond party preferences by considering candidates’ ideological positions and endorsement by union organizations, as well as employees’ support of interest groups. We document considerable within-party variation in the effects on contributions to different candidates. Liberal candidates gain and conservative candidates lose, while moderate candidates are not significantly impacted on average. This suggests that our findings are not only driven by an increased signal of Democratic over Republican party affiliation but reflect shifts between candidates with clearly distinguishable ideological positions. Moreover, we differentiate candidates by whether or not they are supported by unions’ Political Action Committees (PACs). The results indicate that the partisan shift goes beyond union-endorsed candidates, reflecting a more general alignment with Democratic candidates. In addition, we show that our results extend to contributions from employees to PACs: unions mobilize workers to increase donations to union and membership PACs, while managers reduce contributions to corporate PACs. The increased support for labor and civil society interest groups from workers and the reduced support for business interest groups from managers match with the observed ideological realignment in their contributions to candidates. In contrast, we do not find any effects on the donations that the PACs of unionized firms direct to candidates. These results are inconsistent with a potential strategic explanation of changes in donations under which individuals, in particular managers, merely signal goodwill toward the interests of unions.
@@
Third, we shed new light on the spread of political preferences at work by combining establishment-level union election data with an individual-level political outcome. Existing research on the political impact of unions has focused either on individual union members and their households (e.g., Freeman, 2003) or on aggregate outcomes comprising the whole county or state population (e.g., Feigenbaum et al., 2018). By focusing on the unionizing workplace, we are the first to consider within-firm dynamics, particularly the reaction of management –- an out-group that is likely indirectly affected by unionization and a key actor in political influence. Thereby, we expand on studies of politics in the workplace (Colonnelli et al., 2022; Chinoy and Koenen, 2024; Frake et al., 2024), specifically those examining spillovers in political donations between managers and workers (Babenko et al., 2020; Stuckatz, 2022) and the political effects of intergroup contact at work (Andersson and Dehdari, 2021).
@@ 2.1 Unionizing through NLRB Elections
Since 1935, the National Labor Relations Act (NLRA) gives most private-sector workers in the U.S. the right to organize in unions and take collective action, such as bargaining and strikes. Collective bargaining between unions and employers takes place at the establishment level. Traditionally, workers unionize through a secret ballot election at their establishment that is administered by the National Labor Relations Board (NLRB). [^7] While union elections are the primary means by which private-sector workers gain union representation, there are alternative procedures for unionization. First, employers may voluntarily recognize unions without an election through neutrality agreements and “card checks”. These cases are less common, however, since employers generally oppose union organization (Schmitt and Zipperer, 2009). Second, some workers’ bargaining rights are not regulated by the NLRA. For example, the Railway Labor Act determines bargaining rights of airline and railroad workers and several federal, state, and local laws regulate the organization of public-sector employees. The unionization procedure involves three main steps: a petition drive, an election, and certification.
The organizing drive can be initiated either by the workers at an establishment or by a union organization. [^8] See Prummer and Squintani (2023) for an organizational theory on how unionization success can vary between grassroots movements and top-down approaches by established unions. The initiator first needs to gather the signatures of at least 30% of workers in the proposed bargaining unit who thereby express a desire for unionization. With these signatures, an election petition is filed to the NLRB. The NLRB decides whether to accept the petition by ascertaining whether workers in the proposed bargaining unit share common interests that can be adequately represented by the union. If the petition is accepted, the NLRB schedules a secret ballot election, which usually takes place in the workplace. The union wins the election if it obtains a strict majority of the votes cast. In case of union victory, the NLRB certifies the union as the sole authorized representative of employees in the bargaining unit.